OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 6y = –12 10x + 32y = –62 A. infinitely many B. 0 C. 1 D. 2

OpenStudy (mathstudent55):

By looking at the two equations, you can tell they are not one single line, and they are not parallel lines. Automatically, that means they are two intersecting lines. That means 1 solution.

OpenStudy (anonymous):

first, lets reduce the equations.. 2x + 6y = -12 reduces to x + 3y = -6 10x + 32y = -62 reduces to 5x + 12y = - 31 so we have : x + 3y = -6 --->(-5) 5x + 12y = -31 ------------ -5x - 15y = 30 (result of multiplying by -5) 5x + 12y = -31 -------------add -3y = -1 y = 1/3 I am not going to figure x, because I do not need to. There is going to be 1 solution

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

x=-3 y=-1

OpenStudy (anonymous):

(2x+6y=-12)10 (10x+32y=-62)2 20x+60y=-120 (-)20x+64y=-124 -20x-64y=+124 -4y=4 y=-1 2x+6(-1)=-12 2x-6=-12 x=-12+6/2 x=-3 check x=-3 y=-1 in equation 2 10(-3)+32(-1)=-62 -62=-62

OpenStudy (anonymous):

please medal